The work function of a metal is 2.5 × 10^-19 J. (a) Find the threshold frequency for photoelectric emission. (b) If the metal is exposed to a light beam of frequency 6.0 × 10^14 Hz, what will be the stopping potential?

Grade:upto college level

1 Answers

Kevin Nash
9 years ago
Sol. W base 0 = 2.5 * 10^–19 J a) We know W base 0 = hv base 0 v base 0 = W base 0/h = 2.5 * 10^-19/6.63 * 10^-34 = 3.77 * 10^14 Hz = 3.8 * 10^14 Hz b) eV base 0 = hv – W base 0 or, V base 0 = hv – W base 0/e = 6.63 * 10^-34 * 6 *10^14 – 2.5 *10^-19/1.6 *10^-19 = 0.91 V